# Searching for coupled, hyperlight scalars across cosmic history

### v1

@@ III.3 Quasar absorption lines
Figure 3 includes curves that combine all measurements by either marginalizing over each phase $\theta(z_{\mathrm{abs}})$ independently or by fixing $\sin^{2}\theta(z_{\mathrm{abs}})=1/2$ for each absorption redshift $z_{\mathrm{abs}}$. For the varying-$\alpha$ results, the two results are quite similar near the peak, reflecting that the larger collection of $\Delta\alpha$ constraints are effectively randomly scattered about zero. (As explained in Section III.1, the tails of the distribution do fall off as a power law in the phase-marginalized case rather than exponentially.) The difference is much more striking for constraints on the electron mass: while the phase-marginalized result resembles the phase-fixed one near the peaks of the distributions, it exhibits greatly enhanced support at $\Delta m_{e,0}/m_{e,0}\sim-10^{-7}$. This skew arises due to a cluster of higher-redshift observations that themselves skew to negative $\Delta m_{e,0}/m_{e,0}$, which are evident in Fig. 3. When marginalizing over each phase, other measurements centered near zero or positive values of $\Delta m_{e,0}/m_{e,0}$ each only contribute a square-root suppression at negative $\Delta m_{e,0}$ as opposed to an exponential one (when the phases are instead fixed).
Figure 3 displays $1\sigma$ intervals of order $10^{-7}$ for $d_{\lambda}^{(2)}\mathcal{A}_{0}^{2}/2$. As a cross check, we compare these results to those with subsets of the quasar samples partitioned by redshift. For both couplings, selecting only quasars with $z_{\mathrm{abs}}>1.5$ hardly affects the width of the posteriors. However, for this subsample the electron coupling constraints shift further toward negative values (the median lying $1.4\sigma$ rather than $0.5\sigma$ below zero); these observations source the skew evident in Fig. 3. (All other tested subsample exhibit an expected degree of scatter in central values.) Only four of the 14 quasar measurements of $\Delta\mu/\mu$ reside at $z_{\mathrm{abs}}<1.5$, and the bound deriving from them alone degrades by a factor of $\sim 10$. Twelve of the 22 $\Delta\alpha/\alpha$ measurements reside at $z_{\mathrm{abs}}<1.5$, broadening the posterior by a factor of $\sim 2$. Many of the $\Delta\alpha/\alpha$ measurements have $z_{\mathrm{abs}}$ near $1$, with numerous pairs at relatively similar redshifts. Taking subsets that are separated in scale factor by at least $1-10\%$ again only degrades constraints by a factor $\lesssim 2$. Moreover, unlike for the electron coupling, the photon coupling bounds are hardly affected by phase marginalization for any subsample.
@@
In conclusion, we take the result from Fig. 3, $-0.3\lesssim 10^{6}d_{m_{e}}^{(2)}\mathcal{A}_{0}^{2}/2\lesssim 0.07$, as our most conservative estimate of the $95\%$ credible interval. Because this bound is strongly influenced by higher-$z_{\mathrm{abs}}$ observations, it is inappropriate for $m_{\varphi}\lesssim 10^{-31}\leavevmode\nobreak\ \mathrm{eV}$; we therefore take the $z_{\mathrm{abs}}<1$ result, $10^{6}|d_{m_{e}}^{(2)}|\mathcal{A}_{0}^{2}/2\lesssim 5$, to indicate the degree to which bounds degrade at lower masses. (Scalars with masses below $10^{-32}\leavevmode\nobreak\ \mathrm{eV}$ are not captured by our treatment here, as they would not have redshifted sufficiently by the present day for their effect on the present-day values of the constants to be negligible.) Using Eq. 55, the bounds on the quadratic electron coupling is
@@
at $m_{\varphi}\gtrsim 10^{-31}\leavevmode\nobreak\ \mathrm{eV}$, with the bounds for negative and positive couplings respectively degrading by relative factors of $\sim 70$ and $17$ at $m_{\varphi}\approx 10^{-32}\leavevmode\nobreak\ \mathrm{eV}$. For the photon coupling we take $10^{6}|d_{e}^{(2)}|\mathcal{A}_{0}^{2}/2\lesssim 0.6$ as a representative bound. Given that this result is relatively insensitive to the choice of redshift cut (or even whether phases are marginalized over), we take it to apply for all $m_{\varphi}\gtrsim 10^{-32}\leavevmode\nobreak\ \mathrm{eV}$. The resulting bound on the quadratic photon coupling is
@@ III.6 Atomic clocks and pulsar timing arrays
In the regime of hyperlight masses, atomic clocks effectively probe the rate of change of fundamental constants—e.g., a linear drift rate, $\mathrm{d}\ln\lambda/\mathrm{d}t$. Marginalization over the scalar’s unknown phase still penalizes bounds in this case, but with different parametric dependence on $m_{\varphi}$. In general, the theoretical drift rate is $\mathrm{d}g_{\lambda}(\varphi)/\mathrm{d}t=m_{\varphi}d_{\lambda}^{(n)}\sin^{n-1}\vartheta\cos\vartheta/(n-1)!$; per Section III.1, the presence of nodes in phase that scale with $\vartheta^{n-1}$ penalizes bounds by a factor of $m_{\varphi}^{-1}$ for $n=1$ and $2$ and by $m_{\varphi}^{-(n-1)}$ for larger $n$. As such, converting bounds from linear couplings to quadratic couplings requires no additional suppression with $m_{\varphi}$ in the regime where the experimental runtime is much shorter than the scalar’s oscillation period.
The strongest constraint on the photon coupling at low masses to date is that from an optical clock frequency comparison of the electric octupole (E3) and electric quadrupole (E2) transitions of the ${}^{171}\mathrm{Yb}^{+}$ ion with observations spanning over $\sim 6$ years [139], giving
@@
The strongest clock constraint on the electron coupling at low masses is placed by a comparison between a ${}^{171}\mathrm{Yb}$ optical lattice clock and a ${}^{133}\mathrm{Cs}$ fountain microwave clock which yields, over a $\sim 298$-day duration [138],
@@
The bounds listed in this section all only apply for nonclustering scalars with periods well below the total time span covered by the experimental measurements. As discussed in Section III.1, both conditions are satisfied below masses $m_{\varphi}\sim 10^{-28}\leavevmode\nobreak\ \mathrm{eV}$. The ytterbium ion bound is the leading clock-based bound on the quadratic photon coupling and the NANOGrav bound that for the quadratic electron coupling.
@@ III.7 Stars
Ref. [23] considers the production of two scalar quanta in stellar plasmas. Higher dimensional operators like the ones considered here—$\varphi^{2}F_{\mu\nu}F^{\mu\nu}$, $\varphi^{2}\bar{e}e$, and $\varphi^{2}\bar{N}N$ (for $N$ some nucleon)—are suppressed by higher powers of the Planck mass. Thermal production rates therefore depend steeply on temperature, making core-collapse supernovae ($T\sim 30\leavevmode\nobreak\ \mathrm{MeV}$) the most constraining systems. Ref. [23] limits $d_{e}^{(2)}\lesssim 10^{15}$ and $d^{(2)}_{m_{N}}\lesssim 10^{14}$ from photon-pair annihilations ($\gamma+\gamma\rightarrow\varphi+\varphi$) and nucleon-nucleon bremsstrahlung ($N+N\rightarrow N+N+\varphi+\varphi$). Limits on $d_{m_{e}}^{(2)}$ from electron-pair annihilations ($e^{+}+e^{-}\rightarrow\varphi+\varphi$) and Compton-like scattering ($e+\gamma\rightarrow e+\varphi+\varphi$) are further suppressed by $\sim(m_{e}/T)^{2}$ and $\sim\alpha(m_{e}/T)^{2}$, respectively, making the constraints irrelevant both compared to cosmological and laboratory bounds (Fig. 2).
@@
for $d_{M_{\star}}^{(2)}>0$ and a nonrelativistic, nondegenerate star; we expect a suppression to also occur in relativistic and/or degenerate systems. Consequently, background-assisted emission in stellar cores drops off sharply above some (positive) critical coupling, in contrast to signals at the surface of the Earth which still monotonically increase with coupling even beyond the critical value. For a star like the Sun, the condition that the scalar amplitude is well approximated by the cosmological background, Eq. 66, translates to $d^{(2)}_{m_{\odot}}\ll 5\times 10^{5}$. For a neutron star (with radius $\sim 10\leavevmode\nobreak\ \mathrm{km}$ and $1.4$ solar masses), $d^{(2)}_{m_{\mathrm{NS}}}\ll 5$; while this can serve as a first estimate, Eq. 66 cannot be applied rigorously to a very degenerate system like a neutron star. For $d_{M_{\star}}^{(2)}<0$, the solution of Ref. [26] in principle predicts that the scalar diverges at the core for a discrete set of couplings: $\varphi_{\text{core}}(t)\approx\bar{\varphi}(t)\sec\sqrt{|d_{M_{\star}}^{(2)}|GM_{\star}/R_{\star}}$, but the theory becomes nonperturbative in this regime.
@@ IV.1 Big bang nucleosynthesis
A deviation in the value of fundamental constants from their present-day values during big bang nucleosynthesis (BBN) would change the predicted yields of primordial elements, which are currently in excellent agreement with observations [64, 153, 154, 59, 37]. The vast majority of nuclei of primordial origin are either hydrogen or helium-4. Their partitioning is typically quantified by the helium yield by nucleon number, $Y_{p}\equiv 4n_{\mathrm{He}}/(n_{\mathrm{H}}+4n_{\mathrm{He}})$, which differs slightly from the yield by mass $Y_{\mathrm{He}}$ used in Section II.3 and CMB literature. We now review the basic elements of the nucleosynthesis calculations that predict $Y_{p}$ in order to establish its expected scaling with various fundamental parameters, focusing on $\alpha$ and $m_{e}$; we then derive approximate bounds on their early-time values from astrophysical measurements of $Y_{p}$. We assume the parameters are simply fixed to some time-independent values $\alpha_{i}$ and $m_{e,i}$ that differ from their present-day values, $\alpha_{0}$ and $m_{e,0}$, as would be the case for the models we consider in Section II when the scalar field is frozen at early times. Our discussion follows that of Refs. [155, 37].
@@
where $m_{u}$ and $m_{d}$ are the masses of the up and down quarks and $b$ is a proportionality constant. Lattice calculations of the QED contribution to $m_{np}$ yield $m_{np}^{\text{(QED)}}\approx-1\leavevmode\nobreak\ \mathrm{MeV}$ [159, 160], with order $10\%$ precision.666 Calculations by alternative means yield a value $m_{np}^{\text{(QED)}}\approx-0.76\leavevmode\nobreak\ \mathrm{MeV}$ [161] that has often been used in past literature; other recent approaches yield $m_{np}^{\text{(QED)}}\approx-0.58\leavevmode\nobreak\ \mathrm{MeV}$ [162, 163]. See, e.g., Ref. [164] for comparative discussion. Assuming no variation in either $b$ or $\Lambda_{\text{QCD}}$, the neutron-proton mass splitting then varies with $\alpha$ via
@@
The coefficient of $\Delta m_{e,i}$ increases to $0.42$ if the aforementioned effects of shifting $e^{+}$-$e^{-}$ annihilation are neglected, which is consistent with the findings of, e.g., Ref. [153]. For all fundamental constants at their present-day value during BBN, detailed numerical calculations marginalized over Planck’s determination of the baryon abundance [99] yield
@@
$-2.1$ $\lesssim d_{e}^{(2)}\frac{F_{\varphi}}{10^{-2}}\lesssim 1.2$ (94a)
$-16$ $\lesssim d_{m_{e}}^{(2)}\frac{F_{\varphi}}{10^{-2}}\lesssim 9.1.$ (94b)
@@ IV.2 Cosmic microwave background anisotropies
The CMB anisotropies precisely measure the spatial inhomogeneities of the SM plasma at the time of hydrogen recombination as viewed on the sky today. The rates governing recombination dynamics and the diffusion of acoustic waves in the plasma depend on the fine-structure constant and electron mass, providing a probe of their values at redshifts $\gtrsim 1000$. The hyperlight scalar responsible for early-time shifts in $\alpha$ or $m_{e}$ can also reduce the distance to photon last scattering and the rate of structure growth; the former impacts the angular extent of the acoustic peaks on the sky and the latter suppresses the extent to which they are smeared by gravitational lensing. We study these dynamics in detail and perform a quantitative analysis of current cosmological data—measurements of the CMB anisotropies from Planck and low-redshift distances via baryon acoustic oscillations (BAO) and supernova (SN) brightnesses—in Ref. [60].
The CMB alone strongly constrains the early-time value of the fine-structure constant via its modulation of the diffusion rate, which alters the degree to which small-scale anisotropies are suppressed. High-resolution observations like Planck measure the so-called damping tail with sufficient precision to constrain $\alpha_{i}$ at the permille level. The fine-structure constant also affects the rate at which photons and free protons recombine into hydrogen, which impacts the generation of CMB polarization in a manner that scales the overall height of the polarization power spectrum. Planck also measured the $E$-mode component of the polarization spectrum (and its correlation with temperature anisotropies), though less precisely than it did the temperature power spectrum. On its own, Planck yields a $68\%$ posterior interval $\alpha_{i}/\alpha_{0}-1\approx\left(0.7\pm 2.4\right)\times 10^{-3}$, which tightens to $\left(1.3\pm 2.1\right)\times 10^{-3}$ when combined with baryon acoustic oscillation data [60].
@@
The CMB by itself provides weak bounds on electron mass variations (at the $10\%$ level), but BAO and SNe data break degeneracies when combined with CMB data by constraining the late-time expansion history (irrespective of the Hubble tension, i.e., not including the SNe calibrations that lead to the tension). In particular, these low-redshift datasets measure the redshift at which dark energy overtook matter in the Universe’s energy budget; along the CMB’s degeneracy between the electron mass and $\Lambda$CDM parameters, the redshift of matter–dark-energy equality changes substantially. However, measurements of this redshift (or, equivalently, the fraction of Universe’s energy in matter) are mildly discrepant between different probes—in particular, BAO data prefer earlier transitions to dark-energy domination (lower present-day matter fractions) than recent SNe datasets [172, 60, 173]. This discrepancy is puzzling even for standard $\Lambda$CDM cosmology, given that both probes measure cosmological distances over a common interval of the Universe’s history.
Ref. [60] examines the manner in which these differences in low-redshift datasets propagate to constraints on the early-time electron mass (and other $\Lambda$CDM parameters, including the Hubble constant); at the extreme ends, Planck data combined with either SNe data from the Dark Energy Survey [174] or BAO data from the Dark Energy Spectroscopic Instrument [172] yield $10^{2}\left(m_{e,i}/m_{e,0}-1\right)\approx-1.6\pm 1$ and $1.7\pm 0.7$, respectively. Using BAO data from the completed Sloan Digital Sky Survey (SDSS) [175, 176] and Six-degree Field Galaxy Survey (6dFGS) [177] instead yields $0.8\pm 0.7$, while using alternative SNe samples from Pantheon+ [178, 179] and Union3 [180] yield $-0.6\pm 1.1$ and $-2\pm 1.4$, respectively.777 Curiously, the measurement using an older version of the Pantheon sample [181] is $1^{+1.5}_{-1.4}$, a result more consistent with current BAO than current SNe data. This lack of concordance impedes a consensus cosmological constraint, as all dataset combinations yield percent-level measurements that are offset by $1$ to $3\sigma$ at worst. Rather than arbitrarily selecting one measurement to highlight, we present results deriving from each dataset combination in order to be forthright about the discordance.
The preceding discussion made no reference to the gravitational impact of a hyperlight scalar field because current CMB data turn out to disfavor its suppression of structure. While Ref. [60] shows that the late-time increase to the matter density from such a scalar by itself opens up a new parameter degeneracy (that extends not just to the primary CMB anisotropies but BAO and SNe data simultaneously), Planck data do not tolerate the degree of structure suppression induced by a scalar field whose abundance can substantially realize this degeneracy. While this predicted parameter relationship is mildly evident in posterior constraints, its impact on marginal measurements of $m_{e,i}$ or $\alpha_{i}$ is only moderate. Similar conclusions apply to SM neutrinos [60, 173], whose masses are truly free parameters and which have similar phenomenology to hyperlight scalars. The constraints on the abundance of coupled hyperlight scalars from Ref. [60] thus do not differ substantially from prior bounds on uncoupled scalars (at the same mass) [87]: over the mass range $10^{-31}\leavevmode\nobreak\ \mathrm{eV}\lesssim m_{\varphi}\lesssim 10^{-29}\leavevmode\nobreak\ \mathrm{eV}$, the $95$th posterior percentile is $F_{\varphi}\approx 1\%$ to $2\%$. On the other hand, Ref. [173] showed that cosmological neutrino mass measurements also vary substantially among more recent low-redshift datasets; constraints on the abundance of uncoupled, hyperlight scalars are likely also subject to a similar degree of variability.

### v2

@@ III.3 Quasar absorption lines
[…]
@@
[…]
@@
at $m_{\varphi}\gtrsim 10^{-31}\leavevmode\nobreak\ \mathrm{eV}$, with the bounds for negative and positive couplings respectively degrading by relative factors of 70$\sim$ 70∼ 70 and 171717 at $m_{\varphi}\approx 10^{-32}\leavevmode\nobreak\ \mathrm{eV}$. For the photon coupling we take $10^{6}|d_{e}^{(2)}|\mathcal{A}_{0}^{2}/2\lesssim 0.6$ as a representative bound. Given that this result is relatively insensitive to the choice of redshift cut (or even whether phases are marginalized over), we take it to apply for all $m_{\varphi}\gtrsim 10^{-32}\leavevmode\nobreak\ \mathrm{eV}$. The resulting bound on the quadratic photon coupling is
@@ III.6 Atomic clocks and pulsar timing arrays
In the regime of hyperlight masses, atomic clocks effectively probe the rate of change of fundamental constants—e.g., a linear drift rate, $\mathrm{d}\ln\lambda/\mathrm{d}t$. Marginalization over the scalar’s unknown phase still penalizes bounds in this case, but with different parametric dependence on $m_{\varphi}$. In general, the theoretical drift rate is $\mathrm{d}g_{\lambda}(\varphi)/\mathrm{d}t=m_{\varphi}d_{\lambda}^{(n)}\sin^{n-1}\vartheta\cos\vartheta/(n-1)!$; per Section III.1, the presence of nodes in phase that scale with $\vartheta^{n-1}$ penalizes bounds by a factor of $m_{\varphi}^{-1}$ for $n=1$ and 222 and by $m_{\varphi}^{-(n-1)}$ for larger $n$. As such, converting bounds from linear couplings to quadratic couplings requires no additional suppression with $m_{\varphi}$ in the regime where the experimental runtime is much shorter than the scalar’s oscillation period.
The strongest constraint on the photon coupling at low masses to date is that from an optical clock frequency comparison of the electric octupole (E3) and electric quadrupole (E2) transitions of the ${}^{171}\mathrm{Yb}^{+}$ ion with observations spanning over 6$\sim$ 6∼ 6 years [148], giving
@@
The strongest clock constraint on the electron coupling at low masses is placed by a comparison between a ${}^{171}\mathrm{Yb}$ optical lattice clock and a ${}^{133}\mathrm{Cs}$ fountain microwave clock which yields, over a 298$\sim$ 298∼ 298-day duration [147],
@@
The bounds listed in this section all only apply for nonclustering scalars with periods well below the total time span covered by the experimental measurements. As discussed in Section III.1, both conditions are satisfied below masses $m_{\varphi}\sim 10^{-28}\leavevmode\nobreak\ \mathrm{eV}$. The ytterbium ion bound is the leading clock-based bound on the quadratic photon coupling, and the NANOGRAV bound is that for the quadratic electron coupling.
@@ III.7 Stars
[…]
@@
for $d_{M_{\star}}^{(2)}>0$ and a nonrelativistic, nondegenerate star; we expect a suppression to also occur in relativistic and/or degenerate systems. Consequently, background-assisted emission in stellar cores drops off sharply above some (positive) critical coupling, in contrast to signals at the surface of the Earth which still monotonically increase with coupling even beyond the critical value. For a star like the Sun, the condition that the scalar amplitude is well approximated by the cosmological background, Eq. 66, translates to $d^{(2)}_{m_{\odot}}\ll 5\times 10^{5}$. For a neutron star (with radius $\sim 10\leavevmode\nobreak\ \mathrm{km}$ and 1.41.41.4 solar masses), $d^{(2)}_{m_{\mathrm{NS}}}\ll 5$; while this can serve as a first estimate, Eq. 66 cannot be applied rigorously to a very degenerate system like a neutron star. For $d_{M_{\star}}^{(2)}<0$, the solution of Ref. [26] in principle predicts that the scalar diverges at the core for a discrete set of couplings: $\varphi_{\text{core}}(t)\approx\bar{\varphi}(t)\sec\sqrt{|d_{M_{\star}}^{(2)}|GM_{\star}/R_{\star}}$, but the theory becomes nonperturbative in this regime.
@@ IV.1 Big bang nucleosynthesis
A deviation in the value of fundamental constants from their present-day values during BBN would change the predicted yields of primordial elements, which are currently in excellent agreement with observations [64, 162, 163, 59, 37]. The vast majority of nuclei of primordial origin are either hydrogen or helium-4. Their partitioning is typically quantified by the helium yield by nucleon number, $Y_{p}\equiv 4n_{\mathrm{He}}/(n_{\mathrm{H}}+4n_{\mathrm{He}})$, which differs slightly from the yield by mass $Y_{\mathrm{He}}$ used in Section II.3 and CMB literature. We now review the basic elements of the nucleosynthesis calculations that predict $Y_{p}$ in order to establish its expected scaling with various fundamental parameters, focusing on $\alpha$ and $m_{e}$; we then derive approximate bounds on their early-time values from astrophysical measurements of $Y_{p}$. We assume the parameters are simply fixed to some time-independent values $\alpha_{i}$ and $m_{e,i}$ that differ from their present-day values, $\alpha_{0}$ and $m_{e,0}$, as would be the case for the models we consider in Section II when the scalar field is frozen at early times. Our discussion follows that of Refs. [164, 37].
@@
[…]
@@
The coefficient of $\Delta m_{e,i}$ increases to 0.420.420.42 if the aforementioned effects of shifting $e^{+}$-$e^{-}$ annihilation are neglected, which is consistent with the findings of, e.g., Ref. [162]. For all fundamental constants at their present-day value during BBN, detailed numerical calculations marginalized over Planck’s determination of the baryon abundance [108] yield
@@
2.1-2.1- 2.1 $\lesssim d_{e}^{(2)}\frac{F_{\varphi}}{10^{-2}}\lesssim 1.2$ (94a)
16-16- 16 $\lesssim d_{m_{e}}^{(2)}\frac{F_{\varphi}}{10^{-2}}\lesssim 9.1.$ (94b)
@@ IV.2 Cosmic microwave background anisotropies
The CMB anisotropies precisely measure the spatial inhomogeneities of the SM plasma at the time of hydrogen recombination as viewed on the sky today. The rates governing recombination dynamics and the diffusion of acoustic waves in the plasma depend on the fine-structure constant and electron mass, providing a probe of their values at redshifts 1000$\gtrsim$ 1000≳ 1000. The hyperlight scalar responsible for early-time shifts in $\alpha$ or $m_{e}$ can also reduce the distance to photon last scattering and the rate of structure growth; the former impacts the angular extent of the acoustic peaks on the sky and the latter suppresses the extent to which they are smeared by gravitational lensing. We study these dynamics in detail and perform a quantitative analysis of current cosmological data—measurements of the CMB anisotropies from Planck and low-redshift distances via baryon acoustic oscillations (BAO) and supernova (SN) brightnesses—in Ref. [60].
[…]
@@
[…]
Reference [60] examines the manner in which these differences in low-redshift datasets propagate to constraints on the early-time electron mass (and other $\Lambda$CDM parameters, including the Hubble constant); at the extreme ends, Planck data combined with either SNe data from the Dark Energy Survey [183] or BAO data from the Dark Energy Spectroscopic Instrument [181] yield $10^{2}\left(m_{e,i}/m_{e,0}-1\right)\approx-1.6\pm 1$ and 1.7±0.71.7$\pm$ 0.71.7 ± 0.7, respectively. Using BAO data from the completed Sloan Digital Sky Survey (SDSS) [184, 185] and Six-degree Field Galaxy Survey (6dFGS) [186] instead yields 0.8±0.70.8$\pm$ 0.70.8 ± 0.7, while using alternative SNe samples from Pantheon+ [187, 188] and Union3 [189] yield 0.6±1.1-0.6$\pm$ 1.1- 0.6 ± 1.1 and 2±1.4-2$\pm$ 1.4- 2 ± 1.4, respectively.777 Curiously, the measurement using an older version of the Pantheon sample [190] is $1^{+1.5}_{-1.4}$, a result more consistent with current BAO than current SNe data. This lack of concordance impedes a consensus cosmological constraint, as all dataset combinations yield percent-level measurements that are offset by 111 to $3\sigma$ at worst. Rather than arbitrarily selecting one measurement to highlight, we present results deriving from each dataset combination in order to be forthright about the discordance.
[…]